# Shifting AI Efficiency From Model-Centric to Data-Centric Compression

### v2

@@ Abstract
The rapid advancement of large language models (LLMs) and multi-modal LLMs (MLLMs) has historically relied on model-centric scaling through increasing parameter counts from millions to hundreds of billions to drive performance gains. However, as we approach hardware limits on model size, the dominant computational bottleneck has fundamentally shifted to the quadratic cost of self-attention over long token sequences, now driven by ultra-long text contexts, high-resolution images, and extended videos. In this position paper, we argue that the focus of research for efficient AI is shifting from model-centric compression to data-centric compression. We position token compression as the new frontier, which improves AI efficiency via reducing the number of tokens during model training or inference. Through comprehensive analysis, we first examine recent developments in long-context AI across various domains and establish a unified mathematical framework for existing model efficiency strategies, demonstrating why token compression represents a crucial paradigm shift in addressing long-context overhead. Subsequently, we systematically review the research landscape of token compression, analyzing its fundamental benefits and identifying its compelling advantages across diverse scenarios. Furthermore, we provide an in-depth analysis of current challenges in token compression research and outline promising future directions. Ultimately, our work aims to offer a fresh perspective on AI efficiency, synthesize existing research, and catalyze innovative developments to address the challenges that increasing context lengths pose to the AI community’s advancement.
@@ 1 Introduction
The explosive growth of large language models (LLMs) OpenAI:GPT-4 ; Touvron:LLaMA ; Bai:Qwen ; grattafiori2024llama3 ; dong2024internlm ; yang2024Qwen2 ; yang2025Qwen3 ; liu2024deepseek-v3 ; guo2025deepseek-r1 and their multi-modal extensions (MLLMs) liu2023llava ; Liu:LLaVA-1.5 ; Chen:InternVL ; chen2024internvl-2.5 ; zhu2025internvl3 ; wang2024qwen2-vl ; bai2025qwen2.5-vl ; chu2024qwen2-audio over the past few years has driven remarkable gains in AI capabilities. Notably, this unprecedented progress has been largely achieved through increasing model scale across the field, with larger models consistently demonstrating superior performance in reasoning, knowledge acquisition, and task generalization. Indeed, the evolution from early language models with modest parameter counts, such as BERT (117M) devlin2018bert , to today’s state-of-the-art LLMs like Llama 4 llama4 , DeepSeek-R1 guo2025deepseek-r1 , and Qwen-3 yang2025Qwen3 (100B+), demonstrates how each successive model iteration has delivered disproportionate performance improvements through sheer scale. Nevertheless, this relentless pursuit of enhanced performance through increased model size comes at an ever-increasing computational cost. As a result, by early 2024, the dominant source of computational overhead was primarily attributed to the linear growth in parameter count and associated memory requirements.
In response to this scaling trend, the research community has developed numerous model-centric compression techniques. These include model quantization yang2019quantization ; rokh2023comprehensive , network pruning Deep_Compression ; cheng2024survey , knowledge distillation Hinton:KD ; gou2021KD_Survey and low-rank decomposition yu2017low-rank ; idelbayev2020low-rank . These methods directly reduce computational overhead by decreasing model size, and were a natural response to the 2022 to 2024 era, when scaling up model size was the primary driver of performance gains.
As model sizes approach hardware limits, the pace of parameter growth is flattening. Meanwhile, a new computational challenge has emerged: the exponential growth in token sequence lengths. Figure 1 (left) clearly shows that while from 2022 to 2024, model size primarily drove computational costs, reaching around 1000B parameters before stagnating, from 2024 onward, the dominant factor has dramatically shifted to the staggering number of processed tokens, which continues to grow exponentially. This trend is evident across multiple domains: language models now process context lengths orders of magnitude longer than their predecessors llama4 ; yang2025Qwen3 ; liu2024deepseek-v3 , particularly with emerging technologies like long chain-of-thought reasoning guo2025deepseek-r1 and multi-agent systems han2024llm-multi-agent , vision models must handle increasingly high-resolution images li2024llava-ov ; zhu2025internvl3 ; bai2025qwen2.5-vl and longer videos chen2024longvila ; liu2025video-xl-pro , and generative models are tasked with creating higher-resolution images chen2024pixart-sigma ; flux2024 and hour-long videos openai2024sora , requiring substantially more tokens and leading to overwhelming computational overhead. Consequently, by late 2024, the primary computational bottleneck has clearly shifted from model size to the quadratic cost of attention mechanism over these extremely long token sequences.
@@
Figure 1: The evolution of AI efficiency: from model-centric to data-centric compression. From 2022 to 2024, AI model performance gains mainly came from scaling model size, directing efficiency research towards model-centric compression. By mid-2024, with model sizes approaching 1000B parameters, their growth has slowed down. Consequently, the focus has shifted to expanding context length to further enhance model capabilities. This paradigm shift necessitates a transition to data-centric compression, emphasizing context length reduction for model efficiency.
@@
This unprecedented growth in sequence lengths has fundamentally shifted the computational bottleneck from model size to the quadratic cost of attention over long context sequences. Based on this observation, as illustrated in Figure 1 (right), we propose a critical position: the AI community should shift its efficiency optimization paradigm from model-centric to data-centric compression. Specifically, we advocate for token compression - a data-centric compression approach that directly reduces token redundancy in model inputs goyal2020power ; jiang2023llmlingua ; li2024snapkv ; Bolya2023:ToMe ; Bolya2023:ToMeSD ; Chen2024:FastV . Token compression methods address computational overhead by identifying and removing low-information tokens during processing, generally without modifying model architectures or even requiring retraining. Our detailed analysis in Section 3.3 demonstrates that token compression offers compelling advantages in terms of universality, efficiency, and compatibility, positioning it as a promising solution for enabling efficient next-generation LLMs and MLLMs.
@@
  • Evolution of AI Efficiency: We analyze recent developments in long-context AI across various domains, revealing a critical transition from parameter-centric to context-centric computational bottlenecks that necessitates a paradigm shift in efficiency optimization.
  • Unified Formulation of Model Efficiency: We establish a comprehensive mathematical formulation that unifies different perspectives on model efficiency, bridging architectural design, model-centric compression, and data-centric compression approaches through theoretical analysis.
  • Systematic Review of Token Compression: We present a thorough investigation of token compression methods, constructing a unified framework to categorize diverse approaches while analyzing their benefits and trade-offs across different scenarios and tasks.
  • Challenges and Future Directions: We provide an in-depth analysis of current challenges in token compression research and propose promising future directions, aiming to catalyze research efforts toward more efficient and effective compression methods.
@@ 2.1 Token Overhead across Various Domains
The field of artificial intelligence has witnessed remarkable advancements across multiple domains, including natural language processing, computer vision, and content generation. These developments have been largely driven by the introduction of the Transformer architecture vaswani2017:Transformer , which has spawned a wide variety of models. As these domains evolve, we observe a significant increase in token sequence lengths across three main areas:
@@
$\bullet$ Longer Context Length in Language Models Large language models (LLMs) devlin2018bert ; Touvron:LLaMA ; grattafiori2024llama3 ; liu2024deepseek-v3 ; Bai:Qwen ; yang2024Qwen2 ; yang2025Qwen3 ; OpenAI:GPT-4 ; cai2024internlm2 have demonstrated remarkable capabilities in natural language understanding and generation zhao2023survey ; hadi2023survey . The context length these models can handle has expanded dramatically from 2,048 tokens in early models like Llama 1 Touvron:LLaMA to 10M tokens in recent iterations like Llama 4 Scout llama4 . This expansion has led to the emergence of large reasoning models guo2025deepseek-r1 ; hurst2024gpt ; huang2022towards , which focus on complex multi-step problem solving through techniques like long chain-of-thought reasoning liu2024deepseek ; yang2025Qwen3 ; chen2025longcot and multi-agent collaboration talebirad2023Multi-agent-collaboration ; chen2024pandora ; han2024llm-multi-agent .
@@
$\bullet$ Higher Resolution and Longer Video Understanding Building on the success of LLMs, multi-modal large language models (MLLMs) li2022blip ; Liu:LLaVA-1.5 ; Liu:LLaVA-NeXT ; li2024llava-ov ; Bai:Qwen-VL ; Chen:InternVL ; chen2024internvl-2.5 ; zhu2025internvl3 extend these capabilities by integrating vision and text processing wu2023multimodal . The visual inputs these models process have evolved significantly from basic $\times 224$ 224 resolution images in early models like LLaVA liu2023llava to 4K ultra-high-resolution images in InternVL3 zhu2025internvl3 and 8K-frame videos in Video-XL-Pro liu2025video-xl-pro , achieving remarkable performance in tasks involving images bai2025qwen2.5-vl , videos chen2024longvila , and multi-modal reasoning wang2024exploring .
[…]
@@
While these advancements across domains have demonstrated outstanding performance, they now face significant efficiency challenges due to the quadratic cost of attention mechanisms over extremely long token sequences. This growing trend toward longer contexts—whether processing complex reasoning chains in language tasks, high-resolution images and longer videos in understanding tasks, or high-fidelity content in generation tasks—necessitates prioritizing research into model efficiency, particularly in addressing the computational overhead associated with increasing context lengths. Detailed statistical analysis of this trend is presented in the Appendix A.
2.2 Model Efficiency from Different Perspectives
@@
where model efficiency can be optimized from three perspectives: (I) Efficient Computation Architecture aims to design efficient neural architectures $\mathbf{F}$ shen2021-liner-attention ; peng2023rwkv ; gu2023mamba , (II) Model-centric Compression focuses on model weights $\mathbf{W}$ Hinton:KD ; yang2019quantization ; li2017pruning ; yu2017low-rank , and (III) Data-centric Compression targets token sequences from input data $\mathbf{X}$ Rao2021:DynamicViT ; jiang2023llmlingua ; Bolya2023:ToMe ; zou2025:ToCa ; Chen2024:FastV .
(I) Efficient Computation Architecture ($\mathbf{F}$)
@@
Since the computational efficiency of neural networks is determined by their architectural design, optimizing architectures represents a fundamental approach to enhance efficiency. Unlike traditional Transformer architectures with quadratic computational complexity $\mathcal{O}(n^{2})$ in attention vaswani2017:Transformer , where $n$ is sequence length, recent innovations typically achieve linear or sub-quadratic scaling: (i) linear attention reformulates the attention mechanism to achieve linear complexity $\mathcal{O}(n)$, enabling efficient processing while maintaining model capacity katharopoulos2020:transformers-rnn ; shen2021-liner-attention ; (ii) RWKV architecture integrates RNN-like linear scaling $\mathcal{O}(n)$ with transformer-like parallelism for efficient sequence processing peng2023rwkv ; duan2024visionrwkv ; (iii) State Space Models like Mamba leverage structured state space modeling to achieve linear complexity $\mathcal{O}(n)$ while maintaining performance gu2023mamba ; Zhu2024:Vim .While these architectural innovations improve efficiency significantly, they require complete model retraining, motivating the exploration of alternative approaches.
(II) Model-centric Compression ($\mathbf{W}$)
@@
As model parameters directly contribute to computational costs and memory usage, reducing parameter complexity serves as another essential strategy. Model compression can be regarded as model-centric compression, which transforms the original parameter set $\mathbf{W}$ to a reduced set $\mathbf{W}^{\prime}$:
$\mathbf{W}^{\prime}=\boldsymbol{\Gamma}(\mathbf{W}),\quad\text{where}\quad|\mathbf{W}^{\prime}|<|\mathbf{W}|$ (2)
where $\boldsymbol{\Gamma}$ represents the compression operator and ||$\cdot$|| ⋅ | denotes parameter size. Several key approaches have emerged: (i) network pruning removes redundant weights, reducing parameters and computation Network_Pruning ; cheng2024survey ; (ii) model quantization reduces parameter precision from high-bit to low-bit representations yang2019quantization ; rokh2023comprehensive ; (iii) knowledge distillation transfers information from large models to compact ones Hinton:KD ; gou2021KD_Survey ; (iv) low-rank decomposition approximates weight matrices with lower-rank representations yu2017low-rank ; idelbayev2020low-rank . With model sizes plateauing and context lengths growing, research focus has begun shifting from model-level to token-level compression strategies.
(III) Data-centric Compression ($\mathbf{X}$)
Different from model-centric compression, token compression represents a data-centric approach that directly reduces input complexity. Given input data represented as token set $\mathbf{X}$, it produces a reduced representation $\mathbf{X}^{\prime}$:
$\mathbf{X}^{\prime}=\boldsymbol{\Phi}(\mathbf{X}),\quad\text{where}\quad|\mathbf{X}^{\prime}|<|\mathbf{X}|$ (3)
where $\boldsymbol{\Phi}$ is the token compression operator and ||$\cdot$|| ⋅ | denotes token length. This approach complements model compression and has demonstrated significant effectiveness across computer vision Rao2021:DynamicViT ; Bolya2023:ToMe and natural language processing kim2021length ; jiang2023llmlingua .
3 How Token Compression Drives Efficient and Effective Models
@@
In this section, we begin with the research roadmap of data-centric compression (i.e., token compression) in Section 3.1. Then, we comprehensively analyze the benefits of token compression during both training and inference stages in Section 3.2. Finally, we summarize five compelling advantages shared by existing token compression approaches in Section 3.3.
@@
3.1 Research Roadmap - What Makes Token Compression Work?

### v3

@@ Abstract
The advancement of large language models (LLMs) and multi-modal LLMs (MLLMs) has historically relied on scaling model parameters. However, as hardware limits constrain further model growth, the primary computational bottleneck has shifted to the quadratic cost of self-attention over increasingly long sequences by ultra-long text contexts, high-resolution images, and extended videos. In this position paper, we argue that the focus of research for efficient artificial intelligence (AI) is shifting from model-centric compression to data-centric compression. We position data-centric compression as the emerging paradigm, which improves AI efficiency by directly compressing the volume of data processed during model training or inference. To formalize this shift, we establish a unified framework for existing efficiency strategies and demonstrate why it constitutes a crucial paradigm change for long-context AI. We then systematically review the landscape of data-centric compression methods, analyzing their benefits across diverse scenarios. Finally, we outline key challenges and promising future research directions. Our work aims to provide a novel perspective on AI efficiency, synthesize existing efforts, and catalyze innovation to address the challenges posed by ever-increasing context lengths.
Shifting AI Efficiency From
Model-Centric to Data-Centric Compression
[…]
@@ 1 Introduction
Figure 1: The evolution of AI efficiency: from model-centric to data-centric compression. From 2022 to 2024, AI model performance gains were primarily driven by scaling model size, directing efficiency research toward model-centric compression. By 2024, with model sizes approaching 1T parameters, their growth has slowed down. Consequently, the focus has shifted to expanding context length to enhance model capabilities, necessitating a transition to data-centric compression that reduces context length for efficiency.
The explosive growth of large language models (LLMs) OpenAI (2023); Touvron et al. (2023); Grattafiori et al. (2024); Dong et al. (2024); Yang et al. (2025a); Guo et al. (2025a) and their multi-modal extensions (MLLMs) Liu et al. (2024c); Chen et al. (2024d); Zhu et al. (2025); Wang et al. (2024a); Bai et al. (2025) over the past few years has driven remarkable gains in AI capabilities. This progress has been largely achieved by increasing model scale, with larger models consistently showing superior performance in reasoning, knowledge acquisition, and task generalization. The evolution from early models like BERT (117M) Devlin et al. (2018) to today’s state-of-the-art LLMs such as DeepSeek-R1 Guo et al. (2025a) and Qwen-3 Yang et al. (2025a) (100B+) illustrates how scale has delivered substantial performance improvements. Nevertheless, this pursuit of performance through larger models incurs ever-increasing computational costs. As a result, by early 2024, the dominant source of computational overhead was primarily attributed to the linear growth in parameter count and associated memory requirements.
In response to this scaling trend, the research community has developed numerous model-centric compression techniques, including model quantization Yang et al. (2019); Rokh et al. (2023), network pruning Han et al. (2016); Cheng et al. (2024), knowledge distillation Hinton et al. (2015); Gou et al. (2021), and low-rank decomposition Yu et al. (2017); Idelbayev and Carreira-Perpinán (2020). These methods reduce computational overhead by decreasing model size and were a natural response to the 2022–2024 era, when scaling model size was the primary driver of performance gains.
As model sizes approach hardware limits, the pace of parameter growth is flattening. Meanwhile, a new computational challenge has emerged: the exponential growth in context sequence lengths. Figure 1 (left) shows that from 2022 to 2024, model size primarily drove computational costs, reaching around 1T parameters before stagnating. Since then, the dominant factor has shifted dramatically to the staggering number of processed tokens, which continues to grow exponentially. This trend spans multiple domains: language models now handle context lengths orders of magnitude longer than before Meta (2025); Yang et al. (2025a), especially with long chain-of-thought reasoning Guo et al. (2025a) and multi-agent systems Han et al. (2024a); vision models process increasingly high-resolution images Zhu et al. (2025); Bai et al. (2025) and longer videos Qin et al. (2025); and generative models create higher-resolution images Labs (2024) and hour-long videos Brooks et al. (2024), all requiring more tokens and causing substantial computational overhead. Consequently, by late 2024, the primary bottleneck has clearly shifted to the quadratic cost of the attention mechanism over these extremely long context sequences.
@@
This unprecedented growth in sequence lengths has shifted the computational bottleneck from model size to the quadratic cost of attention over long context sequences. Based on this observation, as illustrated in Figure 1 (right), we propose a critical position: the AI community should shift its efficiency optimization paradigm from model-centric to data-centric compression. We advocate for data-centric compression that directly reduces the volume of data processed during model training or inference Jiang et al. (2023); Bolya et al. (2023b); Bolya and Hoffman (2023); Chen et al. (2024b); Lin et al. (2024b). These methods address computational overhead by removing low-information content during processing, typically without modifying model architectures or requiring retraining. Our analysis in Section 3.3 shows that they offer compelling advantages in universality, efficiency, and compatibility, positioning data-centric compression as a promising solution for efficient next-generation LLMs and MLLMs.
@@
  • Evolution of AI Efficiency: We analyze recent developments in long-context AI across various domains, revealing a critical transition from parameter-centric to context-centric computational bottlenecks that necessitate a paradigm shift in efficiency optimization.
  • Unified Formulation of Model Efficiency: We establish a comprehensive mathematical formulation that unifies architectural design, model-centric compression, and data-centric compression within a single expression.
  • Systematic Review of Data-centric Compression: We present a thorough investigation of data-centric compression methods, constructing a unified framework to categorize diverse approaches while analyzing their benefits across different scenarios and tasks.
  • Challenges and Future Directions: We provide an in-depth analysis of current challenges in data-centric compression research and propose promising future directions, aiming to catalyze research efforts toward more efficient and effective compression methods.
@@ 2.1 Token Overhead across Various Domains
The field of AI has witnessed remarkable advancements across multiple domains, including natural language processing, computer vision, and content generation. These developments have been largely driven by the introduction of the Transformer architecture Vaswani et al. (2017), which has spawned a wide variety of models. As these domains evolve, we observe a significant increase in token sequence lengths across three main areas:
@@
(I) Longer Context Length in Language Models: Large language models (LLMs) OpenAI (2023); Touvron et al. (2023); Grattafiori et al. (2024); Liu et al. (2024b); Bai et al. (2023a); Yang et al. (2025a) have demonstrated remarkable capabilities in natural language understanding and generation. The context length LLMs can handle has expanded dramatically from 2,048 tokens in early models like Llama 1 Touvron et al. (2023) to 10M tokens in recent iterations like Llama 4 Scout Meta (2025). This expansion has led to the emergence of large reasoning models Guo et al. (2025a); Yang et al. (2025a), which focus on complex multi-step problem solving through techniques like long chain-of-thought reasoning Liu et al. (2024a) and multi-agent collaboration Han et al. (2024a).
@@
(II) Higher Resolution and Longer Video Understanding: Building on the success of LLMs, multi-modal large language models (MLLMs) Liu et al. (2024c); Li et al. (2024a); Bai et al. (2023b, 2025); Chen et al. (2024d); Zhu et al. (2025); Guo et al. (2025b) extend these capabilities by integrating vision and text processing Wu et al. (2023). Visual inputs processed by MLLMs have evolved from basic $224\times 224$ resolution images in early models like LLaVA Liu et al. (2023) to 4K ultra-high-resolution images in InternVL3 Zhu et al. (2025) and 10K-frame videos in Video-XL-2 Qin et al. (2025), enabling strong performance on image Bai et al. (2025), video Yang et al. (2025b), and multi-modal reasoning tasks Shen et al. (2025).
(III) More Complex Content in Generation Tasks: AI content generation has advanced significantly with the application of Transformers to generative domains Peebles and Xie (2023); Brooks et al. (2024); Li et al. (2024f). Early diffusion models like Stable Diffusion Rombach et al. (2022) generated only $512\times 512$ resolution images. With Transformers now successfully applied to generation Peebles and Xie (2023); Gao et al. (2023); Brooks et al. (2024); Li et al. (2024f), DiT-based models produce high-quality 4K images in PixArt-$\Sigma$ Chen et al. (2024a) and hour-long videos in Sora Brooks et al. (2024). These models capture complex spatiotemporal dependencies, enabling high-fidelity content generation Labs (2024); Yang et al. (2024c); Wan et al. (2025a); Kang et al. (2025).
@@
While these advancements across domains have demonstrated outstanding performance, they now face significant efficiency challenges due to the quadratic cost of attention mechanisms over extremely long token sequences. The growing trend toward longer contexts, whether in complex reasoning chains for language tasks, high-resolution images and longer videos for understanding, or high-fidelity content for generation, necessitates prioritizing research on model efficiency, particularly in mitigating the computational overhead of increasing context lengths. Detailed statistical analysis of this trend is provided in Appendix A.
2.2 AI Efficiency from Different Perspectives
@@
Model efficiency can be optimized from three perspectives: (I) Efficient Computation Architecture designs efficient neural architectures $\mathbf{F}$ Shen et al. (2021); Peng et al. (2023); Gu and Dao (2023), (II) Model-centric Compression reduces model weights $\mathbf{W}$ Hinton et al. (2015); Yang et al. (2019); Li et al. (2017); Yu et al. (2017), and (III) Data-centric Compression compresses token sequences from input data $\mathbf{X}$ Rao et al. (2021); Jiang et al. (2023); Chen et al. (2024b); Zou et al. (2025).
@@
(I) Efficient Computation Architecture ($\mathbf{F}$): Since computational efficiency is determined by architectural design, optimizing $\mathbf{F}$ is fundamental. Unlike Transformers with quadratic attention complexity $\mathcal{O}(n^{2})$ Vaswani et al. (2017), recent methods achieve linear or sub-quadratic scaling: (i) linear attention reformulates attention for $\mathcal{O}(n)$ complexity Katharopoulos et al. (2020); Shen et al. (2021); (ii) RWKV combines RNN-like $\mathcal{O}(n)$ scaling with transformer parallelism Peng et al. (2023); Duan et al. (2024); (iii) State Space Models like Mamba use structured state spaces for $\mathcal{O}(n)$ complexity Gu and Dao (2023); Zhu et al. . These require retraining, motivating alternative approaches.
@@
(II) Model-centric Compression ($\mathbf{W}$): Reducing parameter complexity lowers computational and memory costs. Model compression is model-centric, transforming $\mathbf{W}$ to a smaller $\mathbf{W}^{\prime}$:
$\mathbf{W}^{\prime}=\bm{\Gamma}(\mathbf{W}),\quad\text{where}\quad|\mathbf{W}^{\prime}|<|\mathbf{W}|$ (2)
with $\bm{\Gamma}$ as the compression operator. Key methods include: (i) network pruning Liu et al. (2019); Cheng et al. (2024); (ii) quantization Yang et al. (2019); Rokh et al. (2023); (iii) knowledge distillation Hinton et al. (2015); Gou et al. (2021); (iv) low-rank decomposition Yu et al. (2017); Idelbayev and Carreira-Perpinán (2020). As model sizes plateau and context lengths grow, research is shifting toward data-centric compression.
(III) Data-centric Compression ($\mathbf{X}$): Data-centric compression is a data-centric paradigm that improves efficiency by directly reducing the volume of data processed during training or inference. It encompasses two primary strategies: (i) dataset compression, which selects or distills informative subsets from the training corpus, and (ii) token compression, which directly reduces the length of input sequences during inference. Given an input sequence $\mathbf{X}$, data-centric compression yields a compressed representation $\mathbf{X}^{\prime}$:
$\mathbf{X}^{\prime}=\bm{\Phi}(\mathbf{X}),\quad\text{where}\quad|\mathbf{X}^{\prime}|<|\mathbf{X}|$ (3)
with $\bm{\Phi}$ as the data compression operator. This approach complements model-centric compression and has shown strong effectiveness in vision Rao et al. (2021); Bolya et al. (2023b) and language domains Kim and Cho (2021); Jiang et al. (2023).
3 How Data-centric Compression Drives Efficient and Effective Models
@@
In this section, we begin with the research roadmap of data-centric compression in Section 3.1. Then, we comprehensively analyze the benefits of data-centric compression methods during both training and inference stages in Section 3.2. Finally, we summarize their advantages in Section 3.3.
@@
3.1 Research Roadmap - What Makes Data-centric Compression Work?